OpenStudy (anonymous):

What is the area of a square with a perimeter of 16?

OpenStudy (anonymous):

P = 4s 16 = 4s s = 4 Area = s x s = 4 x 4 = 16 square units

OpenStudy (anonymous):

4*s = 16 s = 16/4 s = 4 a = s^2 a = 4*4 a = 16